# Polarization Diagnostics Applied to Coronal Mass Ejections and the Background Solar Wind

### v1

@@ 3 Polarization Diagnostics of CMEs
Figure 11: pB synthesized twisted croissant model, shown here rotating about the direction of its trajectory to illustrate its 3D structure. MOVIE
@@ 3.1 Twisted Croissant Model
Analytic models are useful for building insight and testing our diagnostics. We employ a croissant type CME model — essentially a hollow shell of density with a twist (Figure 11) The geometry of our model is similar to the graduated cylindrical shell (GCS) model (Thernisien, Howard, and Vourlidas, 2006; Thernisien, Vourlidas, and Howard, 2009), but is created with a different approach which offers the potential to easily create more complex shapes. For example, an overall twist can be applied to the croissant model. The model is initialized by defining the location of the inner tube axis, and parameters such as the twist, orientation, or leg separation are applied to this inner axis. The location of the tube sheath is then calculated in relation to the inner axis by defining the radial distance (or flux tube width), which expands from the CME footpoints to the apex. See Appendix B for further details.
@@
Figure 12: Polarization reveals distance from Thomson Sphere (TS). (left) The CME (twisted croissant model) as in Figure 11 erupting from the West limb, as viewed from the North pole. The lines of sight are shaded in progressively darker shades of blue to illustrate distance from the TS of the points of intersection with the croissant. (right top) $\tau_{COM}$ ground-truth calculated center of mass along the line of sight from the observer (eye shown in left image), and (right bottom) $\tau_{pol}$ calculated from forward-modeled synthetic $pB$ and $tB$. The latter is unsigned, and the negative choice here is appropriate for all points along the edge of the CME as it lies completely behind the TS.
Figure 11 shows synthesized $pB$ for a model CME erupting at the solar West limb. The images afa-f show how the line-of-sight-integrated $pB$ changes if the twisted croissant rotates about the $Y$-axis. Figure 12 shows that the $\tau_{pol}$ calculated from forward-modeled $pB$ and $tB$ reproduces the (unsigned) distance from the TS, $\tau_{COM}$. Note that the lines of sight shown in the left side of the figure each only intersect one point at the edge of the twisted croissant, because the edge that is positive $x$ is also positive $z$ (i.e., above the heliographic equator), while the edge that appears behind the $X$-axis from the North pole view of the left-hand side is below the heliographic equator (negative $z$) as viewed from the Earth and as shown in the images on the right-hand side. Such tangent lines of sight are intrinsically localized as discussed in Section 2.5.
@@
In order to specify the position of the croissant edge in three dimensions, it is useful to calculate it in terms of distance from the $x=0$ plane, i.e., $x_{pol}$. This then can be combined with a projection of the image onto the Observing Plane of Sky ($YZ$; see Figure 1) along with knowledge of how the observers lines of sight intersect said plane, in order to yield XYZX-Y-Z coordinates of the edge (Equations 22-24). Equation 22 has two solutions for $x_{pol}$ corresponding to the $\tau_{pol}$ in front of/behind the TS — we will refer to these as the “front” and “back” solutions. In the case shown in Figure 13, the CME lies entirely behind the TS and so the $x_{pol}$ corresponding to the back solution is the correct one, while the $x_{pol}$ corresponding to the front solution is the so-called “ghost” solution (DeForest, Howard, and Tappin, 2013b), which we now discuss.
Figure 14: Two distinct trajectories arise from time series of polarization data. The solid purple line shows the true trajectory of a CME leading edge as it moves toward the observer, while the dashed purple curve shows a “ghost” trajectory as described in the text. The red, green and blue lines of sight intersect the TS at the positions indicated by colored squares, and the $x=0$ plane as indicated by the colored *. The distance from the TS determined from $PR$, i.e., $\tau_{pol}$, is indicated by orange dots in front of the TS (front solution), and purple circles behind the TS (back solution). Figure adapted from DeForest, Howard, and Tappin (2013b).
@@ 3.2 Tracking CME Trajectories
DeForest, Howard, and Tappin (2013b) explored the front/back of TS ambiguity implicit to analyzing white-light polarization data and demonstrated that the ambiguity can be removed for moving structures in wide-field heliographic images. We briefly review this now, and show how both movies and single images of CME $x_{pol}$ might be used to establish whether a CME is approaching or receding from the observer at Earth.
Figure 15: Approaching CME front as in Figure 14 but with a higher inclination trajectory that crosses the TS closer to the Sun. Note that the first two points along the true trajectory (solid purple line) are solid orange dots (positive $\tau_{pol}$ or front solutions), while the next two points are purple circles (negative $\tau_{pol}$ or back solutions), and the converse is true for the ghost solution (dashed purple line). Figure adapted from DeForest, Howard, and Tappin (2013b).
@@ 3.2.1 Revealing the ghost trajectory
As we have discussed above, the more compact a structure is, the more straightforward it is to locate its position in 3D. An important case is that of the CME leading edge, which by definition has a tangent (compact) intersection with any given line of sight as discussed above (Howard et al., 2013). This is generally true for any surface, not just spherical ones, as it is a consequence of the interior-extremum theorem of calculus. Note that, for most CMEs, the 2D “leading edge” is slightly different than the 3D front propagation location (“nose”) of the CME, an effect that leads to at least some uncertainty in CME direction and propagation speed, even if the CME nose location is fully known. See Malanushenko et al. (2025) for further discussion.
@@ 3.2.2 Imaging halo CMEs
As described above, and discussed in DeForest, Howard, and Tappin (2013b), disambiguation for a CME leading edge is possible with a time series by plotting the radial position vs time of the front and back solutions for the leading edge (Laurent, 2025). But CMEs are more than a single leading edge, and information about their location is encoded in the polarization data from their full extent. This has been demonstrated by Jarolim (2025), who used a physics-informed-neural-network approach to track CME trajectory from synthetic images of eruptions. In this section, we consider in general how images and movies of CME $x_{pol}$ might be used to establish whether a CME is advancing toward or receding from the observer.
@@
Figure 19: Wide vs narrow halo CME approaching the observer. Left column shows wide (top) and narrow (bottom) halo CMEs (see Figure 20) from the observer’s point of view. Right three columns show that despite the fact they have the same projected apparent size and aspect ratio, the polarization ratio yields very different results and immediately shows the wide CME (top row) is farther from the observer, as evidenced by the lower distance from $x=0$ (paler red/blue in images).
Figure 20: Wide vs narrow halo CMEs can be disambiguated by polarization. Red triangles show the correct $x_{pol_{front}}$ solution for the various lines of sight through the narrow CME, while purple triangles show the ghost $x_{pol_{back}}$ solutions on the opposite side of the TS. Similarly, red asterisks show the correct $x_{pol_{front}}$ solutions for the wide CME, while purple asterisks show the ghost $x_{pol_{back}}$ solutions. In this case, only the leading edges for each CME can be thought of as superparticles and located by $x_{pol_{front}}$, while the other lines of sight intersect multiple structures and result in a representation of center of mass equivalent to $x_{COM}$.
@@
In summary: even a single snapshot of a halo CME can yield information about whether it is Earth-approaching or Earth-receding. It can also provide immediate information about whether it is wide and slow vs narrow and fast.
@@ 3.3 Polarization Diagnostics of CME Substructures
CMEs are generally expected to incorporate coherently-twisted magnetic fields, i.e., magnetic flux ropes. The chirality, or handedness of magnetic twist of the flux rope is significant, because they provide information about the direction of the magnetic fields that may interact with the Earth’s magnetic field (Cane, Richardson, and St. Cyr, 2000; Yurchyshyn, Wang, and Abramenko, 2003; DeForest, de Koning, and Elliott, 2017b; Pal, 2022) Figure 21 illustrates this for two examples of right- and left-handed flux ropes expanding out from the Sun.
@@ 3.3.1 Diagnosing chirality
By establishing the location of helical CME substructures using polarization, it is possible to establish chirality from the direction the structure curves when traced from front to back. If clockwise, it is right-handed, while if counterclockwise, it is left-handed (see Figure 22). The ability to do this was demonstrated using coronagraph observations in DeForest, de Koning, and Elliott (2017b), where white light polarization data was used to establish the distance from the $x=0$ plane for a substructure within the erupting CME. Since the CME was observed from two vantage points (SOHO and STEREO), the front vs back solution was easily disambiguated. The $x$ position of the substructure was then traced from front to back along a C-shaped curve, establishing a clockwise direction, and thus right-hand twist. This particular CME was later observed as a magnetic cloud, which also showed right-handed twist, consistent with the determination made from the coronal polarization data.
@@
Polarization analysis of erupting structures such as the prominence modeled here may be done in the low-to-middle corona using the $pB$ and $tB$ data from ESA’s Proba-3 ASPIICS mission (Zhukov et al., 2025). In this region, the TS lies along the $x=0$ plane and does not have the symmetry-breaking curvature found in a wider field of view. The front-back ambiguity is thus harder to disentangle (see, e.g. Dai et al. (2014)). One way around this is to use multiple viewpoints, e.g., polarization data from the STEREO spacecraft (Dai, Wang, and Inhester, 2020). Another way is to monitor an eruption starting in the ASPIICS field of view as it continues into the PUNCH field of view, and use the polarization data from PUNCH to help disambiguate the polarization information in the lower, ASPIICS field of view.
@@
In summary: To locate the position of a CME substructure in 3D, it is first necessary to examine the CME trajectory to eliminate the ghost solution. After this, its 4D space-time trajectory may be obtained and potentially used to diagnose chirality.
@@ 4 Discussion and Conclusions
In this paper, we have demonstrated how the polarization ratio $PR$ contains useful information about the two limits discussed in this paper, i.e., that of a superparticle, or compact mass along the line of sight, and that of spherically symmetric density distribution. The former is dependent on the degree to which the mass is localized, which we have shown has a non-trivial dependence on linear LOS half-width and line-of-sight viewing angle as characterized by the critical width parameter, $\frac{w}{2d}$. The latter represents a local diagnostic of the global density radial falloff that might be used to study variations in the background solar wind as a function of time – for example global responses to large CMEs, or solar cycle timescale variation.
@@
We have used forward models of CMEs to illustrate how the 3d location of the edges of CMEs can be determined and have demonstrated that even with multiple compact masses along a line of sight, the COM density position can be reproduced. We have shown how individual polarization images can establish whether a CME is approaching or receding, and whether it is wide and slow vs narrow and fast. We have demonstrated that while flux rope chirality (or handedness) can be deduced by tracing the curvature of CME substructures, it requires consideration of all projections of the 3D surface mapped out by polarization.
@@
The models used in this paper are idealized, as our goal has been to establish proof-of-concept for the methods described and to deepen our understanding of the sources of uncertainty for polarization diagnostics. We refer the reader to Malanushenko et al. (2025) for an application of these diagnostics to white-light CME data synthesized from more sophisticated global MHD simulations. Of course, real data introduces a range of additional practical difficulties. In particular, imaging data from PUNCH and elsewhere are subject to noise. As described in Section 3.2 of (DeForest et al., 2025), PUNCH has been designed with sensitivity requirements intended to allow 3D tracking of features of expected brightness. There is also the problem of disentangling background structures from CMEs, which time-differencing as employed in Figure 23 cannot fully address. However, we note sophisticated means exist for isolating CME signal from the background without amplifying noise (Morgan, Byrne, and Habbal, 2012), and these have been applied successfully to real time operations (see, e.g., https://solarphysics.aber.ac.uk/Archives/cme˙realtime/realtime˙catalogue.html).
@@
Setting aside noise, we have demonstrated some intrinsic error sources associated with polarization feature localization, which can be quantified if we are in the small-width limit. Just as importantly, we have explored and demonstrated the limits of the superparticle approximation for polarized 3D analysis, including where and how the small-width limit breaks down so that there potential for the information from the polarization to be misinterpreted. As the critical width parameter $\frac{w}{2d}$ transitions toward the infinite width regime, $\tau_{pol}$ loses information about the wedge center position $\tau_{o}$, and the apparent over- and under-estimate in $\tau_{diff}=\tau_{pol}-|\tau_{o}|$ no longer represents an error on position localization, but simply the difference between any given $\tau_{o}$ and the value of $\tau_{pol}$ that happens to characterize that $\frac{w}{2d}$. For this reason, the exact solutions, $\tau_{diff}=0$ (black lines in Figures 8-10), are essentially meaningless and potentially misleading: A broken clock is accurate twice a day.
@@
SIRs are an intermediate case in that they have global scale but form comparatively thin manifolds and hence may be localized along some lines of sight where they are visible. Our assumption of a uniform density wedge is not generally appropriate for SIRs. de Koning et al. (2025) instead consider features of a given angular width $\Delta$ – equivalent to our $\frac{w}{2d}\cos^{2}(\tau_{o})$ in the small-width limit and possessing the same definition of compactness under the assumption of uniform density (i.e., $2\Delta<<1$, equivalent to $\frac{w}{d}\cos^{2}(\tau_{o})<<1$), but find a somewhat more stringent localization requirement when they assume a density profile of $r^{-2}$ within their feature, i.e., $4\Delta<<1$. In between, they examine how $PR$ might be combined with knowledge of either feature angular width or feature central angle to establish the missing variable.
There is thus clear benefit in supplementing polarization diagnostics with information about the size and structure of a feature being studied. Some promising approaches include geometric analyses of feature dimensions (Moraes Filho et al., 2024; López-Portela, Paz Miralles, and Lara, 2025), multiple viewpoints from different spacecraft and/or tomographic reconstructions (Morgan and Cook, 2020; Dai, Wang, and Inhester, 2020; Jackson et al., 2023; Jarolim, 2025), and global MHD simulations (Sachdeva et al., 2021; Lionello et al., 2023; Odstrcil, 2023; Pogorelov et al., 2024; Linan et al., 2025; Provornikova et al., 2025).

### v2

@@ 3 Polarization Diagnostics of CMEs
Figure 11: pB synthesized twisted croissant model, shown here from an observing viewpoint that rotates about the direction of CME trajectory starting from (a) an observer at the heliographic equator with CME on the West limb to (b) $30^{\circ}$, (c) $60^{\circ}$, (d) $90^{\circ}$ (North pole), (e) $120^{\circ}$ (viewing the CME from the solar farside), to (f) $150^{\circ}$. MOVIE
@@ 3.1 Twisted Croissant Model
Analytic models are useful for building insight and testing our diagnostics. We employ a croissant type CME model — essentially a hollow shell of density with a twist (Figure 11). The geometry of our model is similar to the graduated cylindrical shell (GCS) model (Thernisien et al., 2006, 2009), but is created with a different approach which offers the potential to easily create more complex shapes. For example, an overall twist can be applied to the croissant model. The model is initialized by defining the location of the inner tube axis, and parameters such as the twist, orientation, or leg separation are applied to this inner axis. The location of the tube sheath is then calculated in relation to the inner axis by defining the radial distance (or flux tube width), which expands from the CME footpoints to the apex. See Appendix B for further details.
@@
Figure 12: Polarization reveals distance from Thomson Sphere (TS). (left) The CME (twisted croissant model) erupting from the West limb, as viewed from the North pole (the same view shown in Figure 11 d). The lines of sight are shaded in progressively darker shades of blue to illustrate distance from the TS of the points of intersection with the croissant. (right top) $\tau_{COM}$ ground-truth calculated center of mass along the line of sight as seen by the observer at the equator (the same view shown in Figure 11 a), and (right bottom) $\tau_{pol}$ calculated from forward-modeled synthetic $pB$ and $tB$. The latter is unsigned, and the negative choice here is appropriate for all points along the edge of the CME as it lies completely behind the TS.
Figure 11 shows synthesized $pB$ for a model CME erupting at the solar West limb equator. The images afa-f show how the LOS-integrated $pB$ changes if the observer’s viewing angle rotates from equator to pole to equator, or equivalently, if the model twisted croissant rotates about the $Y$-axis. Figure 12 shows that the $\tau_{pol}$ calculated from forward-modeled $pB$ and $tB$ reproduces the (unsigned) distance from the TS, $\tau_{COM}$. Note that the lines of sight shown in the left side of the figure each only intersect one point at the edge of the twisted croissant, because the edge that is positive $x$ is also positive $z$ (i.e., above the heliographic equator), while the edge that appears behind the $X$-axis from the North pole view of the left-hand side is below the heliographic equator (negative $z$) as viewed from the Earth and as shown in the images on the right-hand side. Such tangent lines of sight are intrinsically localized as discussed in Section 2.5. Other lines of sight through the middle of the structure do intersect more than one point of the croissant, and this is explicitly accounted for in the LOS integrals of $pB$ and $tB$. The fact that forward-synthesized $\tau_{pol}$ still reproduces the ground truth $\tau_{com}$, as shown in the right-hand-side of Figure 12, indicates that the polarization ratio can be used in a meaningful way with such superposed structures. We return to this point in Section 3.2.2.
@@
In order to specify the position of the croissant edge in three dimensions, it is useful to calculate it in terms of distance from the $x=0$ plane, i.e., $x_{pol}$. This then can be combined with a projection of the image onto the Observing Plane of Sky ($YZ$; see Figure 1) along with knowledge of how the observers lines of sight intersect said plane, in order to yield XYZX-Y-Z coordinates of the edge (Equations 22-24). Equation 22 has two solutions for $x_{pol}$ corresponding to the $\tau_{pol}$ in front of/behind the TS — we will refer to these as the “front” and “back” solutions. In the case shown in Figure 13, the CME lies entirely behind the TS and so the $x_{pol}$ corresponding to the back solution is the correct one, while the $x_{pol}$ corresponding to the front solution is the so-called “ghost” solution (DeForest et al., 2013b), which we now discuss.
Figure 14: Two distinct trajectories arise from time series of polarization data. The solid purple line shows the true trajectory of a CME leading edge as it moves toward the observer, while the dashed purple curve shows a “ghost” trajectory as described in the text. The red, green and blue lines of sight intersect the TS at the positions indicated by colored squares, and the $x=0$ plane as indicated by the colored *. The distance from the TS determined from $PR$, i.e., $\tau_{pol}$, is indicated by orange dots in front of the TS (front solution), and purple circles behind the TS (back solution). Figure adapted from DeForest et al. (2013b).
@@ 3.2 Tracking CME Trajectories
DeForest et al. (2013b) explored the front/back of TS ambiguity implicit to analyzing white-light polarization data and demonstrated that the ambiguity can be removed for moving structures in wide-field heliographic images. We briefly review this now, and show how both movies and single images of CME $x_{pol}$ might be used to establish whether a CME is approaching or receding from the observer at Earth.
Figure 15: Approaching CME front as in Figure 14 but with a higher inclination trajectory that crosses the TS closer to the Sun. Note that the first two points along the true trajectory (solid purple line) are solid orange dots (positive $\tau_{pol}$ or front solutions), while the next two points are purple circles (negative $\tau_{pol}$ or back solutions), and the converse is true for the ghost solution (dashed purple line). Figure adapted from DeForest et al. (2013b).
@@ 3.2.1 Revealing the ghost trajectory
As we have discussed above, the more compact a structure is, the more straightforward it is to locate its position in 3D. An important case is that of the CME leading edge, which by definition has a tangent (compact) intersection with any given line of sight as discussed above (Howard et al., 2013). This is generally true for any surface, not just spherical ones, as it is a consequence of the interior-extremum theorem of calculus. Note that this 2D tangent to the line of sight (that we refer to as the “leading edge”) is slightly different than the CME front propagation location that is radially farthest from the Sun (sometimes referred to as the CME “nose”; see, e.g., Möstl et al. (2012); Möstl and Davies (2013)) – an effect that leads to an uncertainty in CME direction and propagation speed, even when the CME nose location is fully known (Malanushenko et al., 2025). Note also that our croissant model does not attempt to reproduce the outermost sheath of the CME or associated shock, which are distinct from the magnetic ejecta that drives it and the ICME’s leading edge (Kilpua et al., 2017; Temmer and Bothmer, 2022). It would be well worth doing a full analysis of how the inclusion of polarization localization might improve uncertainties in white-light reconstruction techniques for realistic CME models (e.g., (Verbeke et al., 2023)), but this lies outside the scope of this paper.
@@ 3.2.2 Imaging halo CMEs
As described above, and discussed in DeForest et al. (2013b), disambiguation for a CME leading edge is possible with a time series by plotting the radial position vs. time of the front and back solutions for the leading edge (Laurent, 2025). But CMEs are more than a single leading edge, and information about their location is encoded in the polarization data from their full extent. This has been demonstrated by Jarolim (2025), who used a physics-informed-neural-network approach to track CME trajectory from synthetic images of eruptions. In this section, we consider in general how images and movies of CME $x_{pol}$ might be used to establish whether a CME is advancing toward or receding from the observer.
@@
Figure 19: Wide vs. narrow halo CME approaching the observer. Left column shows wide (top) and narrow (bottom) halo CMEs (see Figure 20) from the observer’s point of view. Right three columns show that despite the fact they have the same projected apparent size and aspect ratio, the polarization ratio yields very different results and immediately shows the wide CME (top row) is farther from the observer, as evidenced by the lower distance from $x=0$ (paler red/blue in images).
Figure 20: Wide vs. narrow halo CMEs can be disambiguated by polarization. Red triangles show the correct $x_{pol_{front}}$ solution for the various lines of sight through the narrow CME, while purple triangles show the ghost $x_{pol_{back}}$ solutions on the opposite side of the TS. Similarly, red asterisks show the correct $x_{pol_{front}}$ solutions for the wide CME, while purple asterisks show the ghost $x_{pol_{back}}$ solutions. In this case, only the leading edges for each CME can be thought of as superparticles and located by $x_{pol_{front}}$, while the other lines of sight intersect multiple structures and result in a representation of center of mass equivalent to $x_{COM}$.
@@
In summary: even a single snapshot of a halo CME can yield information about whether it is Earth-approaching or Earth-receding. It can also provide immediate information about whether it is wide and slow vs. narrow and fast.
@@ 3.3 Polarization Diagnostics of CME Substructures
CMEs are generally expected to incorporate coherently-twisted magnetic fields, i.e., magnetic flux ropes. The chirality, or handedness of magnetic twist of the flux rope is significant, because they provide information about the direction of the magnetic fields that may interact with the Earth’s magnetic field (Cane et al., 2000; Yurchyshyn et al., 2003; DeForest et al., 2017a; Pal, 2022). Figure 21 illustrates this for two examples of right- and left-handed flux ropes expanding out from the Sun.
@@ 3.3.1 Diagnosing chirality
By establishing the location of helical CME substructures using polarization, it is possible to establish chirality from the direction the structure curves when traced from front to back. If clockwise, it is right-handed, while if counterclockwise, it is left-handed (see Figure 22). The ability to do this was demonstrated using coronagraph observations in DeForest et al. (2017a), where white light polarization data was used to establish the distance from the $x=0$ plane for a substructure within the erupting CME. Since the CME was observed from two vantage points (SOHO and STEREO), the front vs. back solution was easily disambiguated. The $x$ position of the substructure was then traced from front to back along a C-shaped curve, establishing a clockwise direction, and thus right-hand twist. This particular CME was later observed as a magnetic cloud, which also showed right-handed twist, consistent with the determination made from the coronal polarization data.
@@
Polarization analysis of erupting structures such as the prominence modeled here may be done in the low-to-middle corona using the $pB$ and $tB$ data from ESA’s Proba-3 ASPIICS mission (Zhukov et al., 2025b, a). In this region, the TS lies along the $x=0$ plane and does not have the symmetry-breaking curvature found in a wider field of view. The front-back ambiguity is thus harder to disentangle (see, e.g. Dai et al. (2014)). One way around this is to use multiple viewpoints, e.g., polarization data from the STEREO spacecraft (Dai et al., 2020). Another way is to monitor an eruption starting in the ASPIICS field of view as it continues into the PUNCH field of view, and use the polarization data from PUNCH to help disambiguate the polarization information in the lower, ASPIICS field of view.
@@
In summary: To locate the position of a CME substructure in 3D, it is first necessary to eliminate the ghost solution. After this, its 4D space-time trajectory may be obtained and potentially used to diagnose chirality.
@@ 4 Discussion and Conclusions
In this paper, we have demonstrated how the polarization ratio $PR$ contains useful information about the two limits discussed in this paper, i.e., that of a superparticle, or compact mass along the line of sight, and that of spherically symmetric density distribution. The former is dependent on the degree to which the mass is localized, which we have shown has a non-trivial dependence on LOS linear half-width and LOS viewing angle as characterized by the critical width parameter, $\frac{w}{2d}$. The latter represents a local diagnostic of the global density radial falloff that might be used to study variations in the background solar wind as a function of time – for example global responses to large CMEs, or solar cycle timescale variation.
@@
We have used forward models of CMEs to illustrate how the 3d location of the edges of CMEs can be determined and have demonstrated that even with multiple compact masses along a line of sight, the COM density position can be reproduced. We have shown how individual polarization images can establish whether a CME is approaching or receding, and whether it is wide and slow vs. narrow and fast. We have demonstrated that while flux rope chirality (or handedness) can be deduced by tracing the curvature of CME substructures, it requires consideration of all projections of the 3D surface mapped out by polarization.
@@
The models used in this paper are idealized, as our goal has been to establish proof-of-concept for the methods described and to deepen our understanding of the sources of uncertainty for polarization diagnostics. We refer the reader to Malanushenko et al. (2025) for an application of these diagnostics to white-light CME data synthesized from more sophisticated global MHD simulations. Of course, real data introduces a range of additional practical difficulties. In particular, imaging data from PUNCH and elsewhere are subject to noise. As described in Section 3.2 of (DeForest et al., 2025), PUNCH has been designed with sensitivity requirements intended to allow 3D tracking of features of expected brightness. There is also the problem of disentangling background structures from CMEs, which time-differencing as employed in Figure 23 cannot fully address. However, we note sophisticated means exist for isolating CME signal from the background without amplifying noise (Morgan et al., 2012), and these have been applied successfully to real time operations (see, e.g., https://solarphysics.aber.ac.uk/Archives/cme˙realtime/realtime˙catalogue.html). Finally, there are the uncertainties intrinsic to 3D reconstruction techniques as referenced above.
@@
Setting aside noise and uncertainty, we have demonstrated some intrinsic error sources associated with polarization feature localization, which can be quantified if we are in the small-width limit. Just as importantly, we have explored and demonstrated the limits of the superparticle approximation for polarized 3D analysis, including where and how the small-width limit breaks down so that there potential for the information from the polarization to be misinterpreted. As the critical width parameter $\frac{w}{2d}$ transitions toward the infinite width regime, $\tau_{pol}$ loses information about the wedge center position $\tau_{o}$, and the apparent over- and under-estimate in $\tau_{diff}=\tau_{pol}-|\tau_{o}|$ no longer represents an error on position localization, but simply the difference between any given $\tau_{o}$ and the value of $\tau_{pol}$ that happens to characterize that $\frac{w}{2d}$. For this reason, the exact solutions, $\tau_{diff}=0$ (black lines in Figures 8-10), are essentially meaningless and potentially misleading: A broken clock is accurate twice a day.
@@
SIRs are an intermediate case in that they have global scale but form comparatively thin manifolds and hence may be localized along some lines of sight where they are visible, for example when the lines of sight are tangent to the stream interface. In general, our assumption of a uniform density wedge is not generally appropriate for SIRs. de Koning et al. (2025) instead consider features of a given angular width $\Delta$ – equivalent to our $\frac{w}{2d}\cos^{2}(\tau_{o})$ in the small-width limit and possessing the same definition of compactness under the assumption of uniform density (i.e., $2\Delta<<1$, equivalent to $\frac{w}{d}\cos^{2}(\tau_{o})<<1$), but find a somewhat more stringent localization requirement when they assume a density profile of $r^{-2}$ within their feature, i.e., $4\Delta<<1$. In between, they examine how $PR$ might be combined with knowledge of either feature angular width or feature central angle to establish the missing variable.
There is thus clear benefit in supplementing polarization diagnostics with information about the size and structure of a feature being studied. Some promising approaches include geometric analyses of feature dimensions (Moraes Filho et al., 2024; López-Portela et al., 2025), multiple viewpoints from different spacecraft and/or tomographic reconstructions (Morgan and Cook, 2020; Dai et al., 2020; Jackson et al., 2023; Jarolim, 2025), and global MHD simulations (Sachdeva et al., 2021; Lionello et al., 2023; Odstrcil, 2023; Pogorelov et al., 2024; Linan et al., 2025; Provornikova et al., 2025).